### v2

@@ 5.2 Evaluating the Precision-Oriented Calibration Algorithm
#Samples RES|VGG Precision Recall #Samples RES|VGG Precision Recall
Airplane 180|177 88%|85% 61%|80% 393|125 80%|81% 68%|86%
Automotive 190|185 91%|82% 68%|73% 290|80 77%|76% 91%|84%
Bird 169|160 89%|87% 79%|78% 441|159 81%|81% 91%|79%
Cat 149|146 91%|83% 41%|53% 547|246 77%|73% 69%|55%
Deer 179|176 90%|85% 82%|91% 526|186 84%|90% 91%|84%
Dog 164|135 88%|80% 50%|68% 465|183 74%|75% 74%|53%
Frog 183|174 88%|83% 100%|99% 452|138 100%|99% 86%|80%
Horse 183|174 85%|84% 64%|79% 404|125 75%|80% 80%|85%
Ship 188|185 90%|84% 95%|94% 260|64 94%|93% 90%|83%
Truck 185|184 89%|87% 91%|91% 331|100 91%|91% 88%|86%
An analysis of these results reveals that DEM accomplishes a 100% detection rate within the Frog category — a level which could be deemed sufficient for the highest aerospace certification level. Other categories achieved lower results, that might be sufficient for lower certification levels. Furthermore, Table 2 illustrates that employing the amalgamation of Algorithm 1 and Algorithm 2 affords the most favorable perspective on sample robustness.
6 Discussion and Future Work
We presented here the DEM tool, which uses a novel, efficient probabilistic certification approach to determine the trustworthiness of individual DNN predictions. The proposed technique has several advantages: (i) it is applicable to black-box DNNs; (ii) it is computationally efficient during inference, using pre-calibrated parameters; and (iii) our favorable evaluation results highlight DEM’s success in identifying adversarially distorted inputs. DEM could enable the selective use of DNN outputs when they are safe, while requesting human oversight when needed.
Moving forward, we plan to extend the study to cover more specific aviation use-cases and related DNNs. In the longer term, we hope DEM could assist in creating certified DNN co-pilots. We hope that including DEM variance profiles within functional hazard analyses may facilitate regulatory acceptance, allowing the potential of AI to be realized in the aerospace domain.

### v3

@@ V-B Evaluating the Precision-Oriented Calibration Algorithm
#Samples RES, VGG Precision Recall #Samples RES, VGG Precision Recall
Airplane 180,177 88%,85% 61%,80% 393,125 80%,81% 68%,86%
Automotive 190,185 91%,82% 68%,73% 290,80 77%,76% 91%,84%
Bird 169,160 89%,87% 79%,78% 441,159 81%,81% 91%,79%
Cat 149,146 91%,83% 41%,53% 547,246 77%,73% 69%,55%
Deer 179,176 90%,85% 82%,91% 526,186 84%,90% 91%,84%
Dog 164,135 88%,80% 50%,68% 465,183 74%,75% 74%,53%
Frog 183,174 88%,83% 100%,99% 452,138 100%,99% 86%,80%
Horse 183,174 85%,84% 64%,79% 404,125 75%,80% 80%,85%
Ship 188,185 90%,84% 95%,94% 260,64 94%,93% 90%,83%
Truck 185,184 89%,87% 91%,91% 331,100 91%,91% 88%,86%
Table V-B highlights the fact that taking into account both Precision and Recall affords a more favorable analysis of sample robustness.
VI Discussion and Future Work
We presented here the DEM tool, which uses a novel, efficient probabilistic certification approach to determine the trustworthiness of individual DNN predictions. The proposed technique has several advantages: (i) it is applicable to black-box DNNs; (ii) it is computationally efficient during inference, using pre-calibrated parameters; and (iii) our evaluation results highlight DEM’s success in identifying adversarially distorted inputs. DEM could enable the selective use of DNN outputs when they are safe, while requesting human oversight when needed.
Moving forward, we plan to extend the study to cover more specific aviation use cases and related DNNs. In the long run, we hope that DEM will aid in creating certified DNN co-pilots. We further hope that including the DEM variance profile within functional hazard analyses may facilitate regulatory acceptance of classifier DNNs, allowing us to realize more of the AI potential in the aerospace domain.